OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x = 6 x = 4 x = -3 x = -2

OpenStudy (anonymous):

the midpoint has coordinates:\[(\frac{ x _{1} + x _{2} }{ 2 }, \frac{ y _{1} + y _{2} }{ 2 })\] So, \[2 = \frac{ x - 2 }{ 2 }\]

OpenStudy (anonymous):

So, now just solve for x

OpenStudy (anonymous):

To solve for x, multiply both sides of that last equation by 2. Then add 2 to both sides.

OpenStudy (anonymous):

@tcarroll010 I got 6 on the left side

OpenStudy (anonymous):

Very good! Good job!

OpenStudy (anonymous):

Awesome!! Thank you very much :)! @tcarroll010

OpenStudy (anonymous):

You're welcome!